Which equation could be used to solve this problem?

The sum of two consecutive odd integers is 124. Find the integers

Answer by stanbon(75887)

The sum of two consecutive odd integers is 124. Find the integers
----------
1st: 2x-1
2nd: 2x+1
---------------------
Equation:
2x-1 + 2x+1 = 124
4x = 124
x = 31
-----
1st: 2x-1 = 61
2nd: 2x+1 = 63
-------------------
Cheers,
Stan H.
===================